OpenStudy (anonymous):

how do I square a binomial (2a-2b)^2??? Can someone show me formula?

OpenStudy (anonymous):

hello rosanne!

OpenStudy (anonymous):

hi satellite!!! how r u? I have a final on mon for this class. Im going to have to bust my retricegetting this down perfect.

OpenStudy (anonymous):

bust my buttt*

OpenStudy (anonymous):

yeah make sure to bust your retrice. whatever that is! lol

OpenStudy (aroub):

yes sure! all u have to do is the follow the rule... perfect squares:(x-b)^2= x^2 -2xb+b^2

OpenStudy (anonymous):

just make sure when you see \[(\text{this + that})^2\] to write \[(\text{this + that})\times (\text{this + that})\] and you will get it right

OpenStudy (anonymous):

\[(2a-2b)^2=(2a-2b)(2a-2b)=4a^2-4ab-4ab+4b^2\] \[=4a^2-8ab+4b^2\]

OpenStudy (anonymous):

why did u subtract?

OpenStudy (anonymous):

be careful with the plus and minus signs, and make sure to do FOUR multiplications and not say something silly like \[(a+b)^2=a^2+b^2\] which is absolutely wrong. math teacher hate that

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

\[(2a-2b)^2=(2a-2b)(2a-2b)\] so far so good?

OpenStudy (anonymous):

why r u subtracting?

OpenStudy (anonymous):

i just copied down exactly what you wrote. i have done nothing else. you wrote \[(2a-2b)^2\] yes?

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok we continue

OpenStudy (anonymous):

you have to do 4 multiplications.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first step is?

OpenStudy (anonymous):

\[2a\times 2a\] \[2a\times (-2b)\] \[(-2b)\times 2a\] \[(-2b)\times (-2b)\] sometimes this is called foil for "first outer inner last"

OpenStudy (anonymous):

that is after you write \[(2a-2b)^2=(2a-2b)(2a-2b)\] then you do 4 multiplications as i wrote above

OpenStudy (anonymous):

now we compute as follows \[(2a)\times (2a)=4a^2\] \[2a\times (-2b)=-4ab\] \[(-2b)\times (2a)=-4ab\] \[(-2b)\times (-2b)=+4b^2\] are these clear?

OpenStudy (anonymous):

wait u lost me why did u put a -2 '

OpenStudy (anonymous):

because i am thinking as follows \[(2a-2b)=2a+(-2b)\] the minus sign comes with the variable. you cannot ignore it

OpenStudy (anonymous):

so i would say the first term in \[2a-2b\] is \[2a\] and the second term is \[-2b\] you have to keep the minus sign

OpenStudy (anonymous):

oh ok i c

OpenStudy (anonymous):

ok good. so now we have the following after doing FOUR multiplications: \[4a^2-4ab-4ab+4b^2\] remember the signs count. they are important

OpenStudy (anonymous):

make sure you know that the signs are correct. now the last step is to comine \[-4ab-4ab\] and that is \[-8ab\]

OpenStudy (anonymous):

because you think "minus 4 minus 4 is minus 8"

OpenStudy (anonymous):

so your "final answer" is \[4a^2-8ab+4b^2\]

OpenStudy (anonymous):

we can do another one if you like

OpenStudy (anonymous):

I still like just standard distribution: \[(2a - 2b)^2 \]\[= (2a - 2b)(2a-2b)\]\[=2a(2a-2b) - 2b(2a-2b)\]\[=4a^2 - 4ab - 4ab +4b^2\]\[=4a^2 - 8ab + 4b^2\]

OpenStudy (anonymous):

It always works with any number of factors and terms.